jennylove:

3. Light coming from a fish makes an incidence angle of 30° to normal under the water. The index of refraction of water is 1.33. A fisherman is looking at the fish through air. At what angle with the normal will the fish appear to the fisherman? Use Snell’s law: .

jennylove:

im unsure how to solve this

justjm:

You know the formula for snell's law? \(n_{1}\sin\theta_{1}=n_{2}\sin\theta_{2}\) where n = refractive and incident indices \(\theta\) = refractive and incident angle Now it seems like the light coming from a fish makes an incident angle of 30°, so \(\theta_1=30°\). The refractive index of water is 1.33, so \(n_1=1.33\). I think where you are confused is that you probably did not realize that the light wave of the image underwater leaves the water medium and enters air, and the refractive index for air is simply 1. Hence \(n_2=1.00\). Now you can find \(\theta_2\)

jennylove:

oh okay so it would end up looking like tihs 

jennylove:

yes i forgot about it being under water! that makes sense

jennylove:

if i gave you my end result could you check if im right?

justjm:

sure

jennylove:

i got 41.68 degrees , is this correct?

justjm:

Yes correct and correct sig figs too (:

jennylove:

Awesome! thank you so much

justjm:

Np